OpenStudy (jaceykilla):

Dave has a three-year college loan for $10,000. The lender charges a simple interest rate of 6 percent. How much interest will he have to pay?

OpenStudy (koffeemonster):

Hello, Jacey! This would be a mathematic question, but I will help anyway! c: ------- Provided information: Principal: $10,000 Rate: 0.06 Time: 3 years ------- SI = P * R * T SI = 10,000 * 0.06 * 3 10,000 multiplied by 0.06 equals 600. 600 multiplied by 3 equals 1,800. Answer: $1,800. ------- Alternative: Months in total of three years: 36 Principal: $10,000 Rate: 6% $10,000 Principal at 6% annual interest rate is equivalent to a monthly interest payment of $50. $50*36 50 multiplied by 36 equals 1,800. Answer: $1,800. ------- The amount of interest Dave would have to pay would be a total of $1,800.